PSQ is a focal chord of ellipse 4x^2+9y^2=36 such that SP=4. if S’ is the other focus then write the value of S’Q
 

Grade:12

1 Answers

Gorav Gupta
6 years ago
Hi Tanisha,
 
The semi latus rectum is the Harmonic mean of SP and SQ.
i.e. (b^2)/a = (2 * SP * SQ)/(SP + SQ)
Get the value of SQ from here.
 
Now, S’Q + SQ = 2 * Major axis = 2 * a 
Where a = 3
The answer will be 26/5